Ticket Number: HEA-NOC/20080127-5 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: HEA-NOC
Ticket Opened: 20080127 16:02 UTC Problem Start: 20080127 14:36 UTC
Ticket Closed: 20080201 16:17 UTC Problem End: 20080131 08:15 UTC
Site/line: St. Angela's
There was a brief flap to St. Angela's (~4 seconds, only observed on
Traffic to St. Angela's.
20080127 15:22 CS Reported to BT, ref: 1-551828715
20080127 17:53 CS BT report seeing no alarms. Will monitor for another
20080128 15:38 LT No further issues for 25 hours. Closing ticket.
20080131 09:53 LT This link flapped several times again this morning,
between 08:09 and 08:15. There were 4 flaps lasting 10 seconds or less
(link did not report as down during these), and one link outage lasting
2.5 minutes. Contacting BT.
20080131 10:16 LT BT reference: 1-553173615.
20080131 16:46 LT BT have only one record of an event on this link, at
07h36 (which does not coincide with the longest outage this morning and is
11 minutes later than the nearest recorded flap on our kit). They will dig
further and get back to us tomorrow.
20080201 16:14 LT BT have determined that the cause of the outages was
environmental issues affecting a radio leg of the link. Other radio links
in the area were similarly affected at the time and it is considered
particularly a risk at this time of year.
The link has been stable since. Closing the ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080127-5
HEAnet Limited. Registered in Ireland, No. 275301.